# VideoPASTA [Uncaptioned image]: 7K Preference Pairs That Matter for Video-LLM Alignment

### v1

@@ 3.4 Preference Data Filtering
We generate three adversarial examples for each failure mode (spatial, temporal, cross-frame) per query and use Qwen2.5-32B [61] as a lightweight verification step to ensure the adversarial examples are genuinely incorrect. We prompt Qwen2.5-32B with a textual comparison task to verify that each adversarial example introduces a deliberate misalignment rather than simply rephrasing the correct answer (prompt provided in Appendix, Figure 13). Adversaries that are too similar to the aligned samples or lack clear contradictions are discarded and regenerated. Similarly, we perform a “sanity check” on preferred responses to ensure they correctly align with the queries without errors. This filtering process creates preference pairs that accurately represent the targeted failure modes, enabling more precise alignment during DPO.
@@ 3.5 Training Process
VideoPASTA leverages structured preference pairs to address distinct failure modes in video understanding through DPO. We begin by partitioning the preference dataset $\mathcal{D}=\{(V,q,r^{+},r^{-})\}$ into three subsets: $\mathcal{D}_{s}$, $\mathcal{D}_{t}$, and $\mathcal{D}_{c}$, corresponding to spatial, temporal, and cross-frame alignment, respectively. Each data point in the preference dataset is a tuple $(V,q,r^{+},r^{-})$, where $V$ represents the input video, $q$ denotes the query regarding the video content, $r^{+}$ is the preferred response, and $r^{-}$ is the adversarial response.
For a video-language model $M_{\theta}$, we define the DPO loss for a single preference pair as:
$\Delta(V,q,r^{+},r^{-})$ $=\log p_{\theta}(r^{+}\mid V,q)-\log p_{\theta}(r^{-}\mid V,q),$ (2)
@@
where $\sigma$ is the sigmoid function, $\lambda$ is a scaling factor, and the term inside the brackets represents the log probability difference between the preferred and adversarial responses. We then compute the DPO loss separately for each subset of preference pairs and weight them by $\alpha$, $\beta$, and $\gamma$ to reflect the relative importance of spatial, temporal, and cross-frame alignment. The overall training objective is:
$\mathcal{L}=\alpha\,\mathbb{E}_{\mathcal{D}_{s}}\big{[}\mathcal{L}_{\text{DPO}}\big{]}+\beta\,\mathbb{E}_{\mathcal{D}_{t}}\big{[}\mathcal{L}_{\text{DPO}}\big{]}+\gamma\,\mathbb{E}_{\mathcal{D}_{c}}\big{[}\mathcal{L}_{\text{DPO}}\big{]}.$ (3)
@@
Model TempCompass PerceptionTest NeXTQA MVBench Vinoground MLVU LongVideoBench VideoMME (1) Baseline Models Qwen2.5-VL [3] 71.7 68.6 75.8 65.2 27.9 68.7 60.7 62.2 + SFT 71.8 69.1 77.2 65.5 28.0 68.8 60.9 62.5 + Hound-DPO [70] 70.3 67.6 76.1 65.7 27.8 66.4 56.3 63.2 + TPO [28] 71.5 69.0 77.6 65.3 28.1 68.9 59.2 64.2 (2) State-of-the-Art Models VideoLLaMA2 [8] 43.4 51.4 - 54.6 22.1 35.5 - 47.9 Kangaroo [32] - - - 61.0 61.0 54.8 56.0 LLaVA-NeXT-Video [71] 53.0 48.8 53.5 53.1 17.8 - 49.1 46.5 LongVA [69] - - - - 58.8 51.3 52.6 LLaVA-NeXT-Interleave [26] 54.1 51.2 67.0 46.5 15.2 52.5 44.8 48.3 InternVL2.5 [7] 68.3 62.2 77.0 69.8 29.4 59.5 52.9 57.9 Qwen2-VL [52] 68.9 62.3 75.7 64.9 24.8 57.5 55.6 55.3 LLaVA-OneVision [25] 64.5 57.1 79.3 56.7 25.0 64.9 56.3 58.2 LLaVA-Video [73] 66.4 67.9 74.2 58.6 24.3 66.5 58.2 62.4 (3) Preference-Optimized Models LLaVA-Hound-DPO [70] 55.5 45.1 61.6 36.6 23.8 41.1 36.7 34.2 i-SRT [1] 56.0 47.0 63.0 36.3 23.7 39.9 38.2 34.7 LLaVA-Video-TPO [28] 66.6 66.3 77.8 56.7 24.0 66.3 58.3 62.4 VideoPASTA [Uncaptioned image] 72.3+0.84% 69.4+1.17% 77.3+1.97% 66.3+1.69% 28.3+1.43% 69.2+0.73% 61.5+1.31% 64.1+3.05%
Table 1: Comprehensive evaluation of VideoPASTA against leading video understanding models. We compare VideoPASTA with (1) baseline models built on Qwen2.5-VL, (2) current state-of-the-art models, and (3) models enhanced through preference optimization. The best scores are in bold, and the second-best scores are underlined. For VideoPASTA, relative performance improvements over the Qwen2.5-VL foundation model are indicated as subscripts. All results, except those marked with †, are reproduced using LMMs-Eval [67].
@@ 4 Experiments and Evaluation
For training, we sample 3,000 videos from ActivityNet [64]. Our structured adversarial sampling pipeline produces  $90,000$ preference pairs which after filtering are reduced to $7,020$ (additional details included in Appendix §A.3). We fine-tune Qwen2.5-VL [3] with the SWIFT [75] framework for efficient adaptation. Training and evaluation are performed on two NVIDIA L40S GPUs (48GB), with a maximum frame limit of 32 to avoid CUDA OOM errors. We use LoRA [17] with $\alpha=\beta=8$ and a DPO scaling factor of $\beta=0.1$. The model converges in one epoch, requiring just 6 hours of training. Evaluation is done using LMMs-Eval [67] for a fair comparison to prior work. Additional experiments on data quality (§A.2), preference learning efficiency (§A.2.2), dataset samples (§A.3.2), qualitative samples (§A.4) and prompt design (§A.5) are provided in appendix.
We evaluate on general video understanding benchmarks: TempCompass [33] (temporal understanding), PerceptionTest [37] (visual perception), NeXTQA [58] (compositional reasoning), MVBench [27] (multi-task reasoning), and Vinoground [66] (dense temporal reasoning). For long-form evaluation, we use LongVideoBench [56] (hour-long videos), MLVU [76] (multi-task, 3-minute to 2-hour videos), and VideoMME [13] (6 visual domains, 30 subfields, 11-second to 1 hour videos).

### v3

@@ 3.4 Preference Data Filtering
We generate three adversarial examples for each failure mode per query and use an open-source LLM, Qwen2.5-32B Yang et al. (2024) as a lightweight verification step to ensure the adversarial examples are genuinely incorrect. We prompt Qwen2.5-32B with a textual comparison task to verify that each adversarial example introduces a deliberate misalignment rather than simply rephrasing the correct answer. Adversaries that are too similar to the aligned samples or lack clear contradictions are discarded and regenerated. Similarly, we perform a “sanity check” on aligned responses to ensure they correctly align with the queries without errors. This filtering process creates preference pairs that accurately represent the targeted failure modes, enabling more precise alignment during DPO.
Model TempCompass (Avg.) PerceptionTest (val_mc) NeXTQA (mc_test) MVBench MLVU (dev) LongVideoBench (val_v) VideoMME (w/o sub) State-of-the-Art Models VideoLLaMA2 (Cheng et al., 2024) 43.4 51.4 - 54.6 35.5 - 47.9 Kangaroo Liu et al. (2024a) - - - 61.0 61.0 54.8 56.0 LLaVA-NeXT-Video (Zhang et al., 2024c) 53.0 48.8 53.5 53.1 - 49.1 46.5 LongVA Zhang et al. (2024b) - - - - 58.8 51.3 52.6 Qwen2-VL (Wang et al., 2024b) 68.9 62.3 75.7 64.9 57.5 55.6 55.3 LLaVA-Video (Zhang et al., 2024d) 66.4 67.9 74.2 58.6 66.5 58.2 62.4 Off-the-Shelf Preference-Optimized Models LLaVA-Hound-DPO (Zhang et al., 2025) 55.5 45.1 61.6 36.6 41.1 36.7 34.2 i-SRT Ahn et al. (2024a) 56.0 47.0 63.0 36.3 39.9 38.2 34.7 LLaVA-Video-TPO (Li et al., 2025b) 66.6 66.3 77.8 56.7 66.3 58.3 62.4 Model-Agnostic Preference Optimization using VideoPASTA LLaVA-NeXT-Interleave (Baseline) 54.1 51.2 67.0 46.5 52.5 44.8 48.3      + SFT 54.3 51.5 67.4 46.8 52.7 45.0 48.5      + Hound-DPO (Zhang et al., 2025) 51.7 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-2.4)}$ 49.5 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.7)}$ 65.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.2)}$ 44.3 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-2.2)}$ 50.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-2.3)}$ 42.5 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-2.3)}$ 46.7 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.6)}$      + TPO (Li et al., 2025b) 54.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.2)}$ 52.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.2)}$ 68.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.5)}$ 47.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.4)}$ 53.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.1)}$ 46.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.7)}$ 49.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.3)}$      + VideoPASTA[Uncaptioned image] 56.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.3)}$ 53.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.6)}$ 70.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.1)}$ 49.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.5)}$ 55.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.3)}$ 47.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.1)}$ 51.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.1)}$ LLaVA-OneVision (Baseline) 64.5 57.1 79.3 56.7 64.9 56.3 58.2      + SFT 64.6 57.4 79.3 56.9 65.1 56.5 58.1      + Hound-DPO (Zhang et al., 2025) 63.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.3)}$ 55.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.3)}$ 78.1 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.2)}$ 55.3 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.4)}$ 63.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.7)}$ 54.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.5)}$ 56.9 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.3)}$      + TPO (Li et al., 2025b) 65.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.1)}$ 58.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.3)}$ 80.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.3)}$ 57.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.2)}$ 65.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.9)}$ 57.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.2)}$ 59.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.0)}$      + VideoPASTA[Uncaptioned image] 67.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.7)}$ 60.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.2)}$ 81.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.5)}$ 59.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.4)}$ 67.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.6)}$ 58.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.2)}$ 60.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.9)}$ InternVL2.5 (Baseline) 68.3 62.2 77.0 69.8 59.5 52.9 57.9      + SFT 68.2 62.3 77.4 70.4 59.4 53.0 58.1      + Hound-DPO (Zhang et al., 2025) 66.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.5)}$ 61.0 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.2)}$ 74.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-2.2)}$ 64.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-5.6)}$ 60.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.7)}$ 54.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.4)}$ 54.6 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-3.3)}$      + TPO (Li et al., 2025b) 68.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-0.1)}$ 62.0 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-0.2)}$ 77.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.2)}$ 68.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.0)}$ 61.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.0)}$ 58.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+5.2)}$ 60.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.1)}$      + VideoPASTA[Uncaptioned image] 71.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.6)}$ 66.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.9)}$ 80.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.7)}$ 73.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+4.0)}$ 63.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+3.9)}$ 58.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+5.2)}$ 62.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+4.1)}$ Qwen2.5-VL (Baseline) 71.7 68.6 75.8 65.2 68.7 60.7 62.2      + SFT 71.8 69.1 77.2 65.5 68.8 60.9 62.5      + Hound-DPO (Zhang et al., 2025) 70.3 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.4)}$ 67.6 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.0)}$ 76.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.3)}$ 65.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.5)}$ 66.4 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-2.3)}$ 56.3 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-4.4)}$ 63.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.0)}$      + TPO (Li et al., 2025b) 71.5 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-0.2)}$ 69.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.4)}$ 77.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.8)}$ 65.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.1)}$ 68.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.2)}$ 59.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.5)}$ 64.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+2.0)}$      + VideoPASTA  [Uncaptioned image] 72.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.6)}$ 69.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.8)}$ 77.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.5)}$ 66.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.1)}$ 69.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.5)}$ 61.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+0.8)}$ 64.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(+1.9)}$
Table 1: Comprehensive evaluation of VideoPASTA against leading (7B) video understanding models. The best scores are in bold, and the second-best scores are underlined. Results marked with † are from the original papers; all other results are reproduced using LMMs-Eval Zhang et al. (2024a).
@@ 3.5 Training Process
VideoPASTA leverages structured preference pairs to address distinct failure modes in video understanding through DPO. We begin by partitioning the preference dataset $\mathcal{D}=\{(V,q,r^{+},r^{-})\}$ into three subsets: $\mathcal{D}_{s}$, $\mathcal{D}_{t}$, and $\mathcal{D}_{c}$, corresponding to spatial, temporal, and cross-frame alignment, respectively. For a video-language model $M_{\theta}$, we define the DPO loss for a single preference pair as:
$\Delta(V,q,r^{+},r^{-})={}$ $\log p_{\theta}(r^{+}\mid V,q)$ (1)
$-\log p_{\theta}(r^{-}\mid V,q),$
@@
where $\sigma$ is the sigmoid function and $\lambda$ is a scaling factor. We then compute the DPO loss for each subset of preference pairs, weighted by $\alpha$, $\beta$, and $\gamma$ for spatial, temporal, and cross-frame alignment, respectively. The overall training objective is:
$\begin{split}\mathcal{L}={}&\alpha\,\mathbb{E}_{\mathcal{D}_{s}}\big[\mathcal{L}_{\text{DPO}}\big]+\beta\,\mathbb{E}_{\mathcal{D}_{t}}\big[\mathcal{L}_{\text{DPO}}\big]\\ &+\gamma\,\mathbb{E}_{\mathcal{D}_{c}}\big[\mathcal{L}_{\text{DPO}}\big].\end{split}$ (2)
@@ 4 Experiments and Evaluation
Figure 3: Qualitative comparison of VideoPASTA against Qwen2.5-VL Bai et al. (2025). Examples show VideoPASTA improves (1) Spatial reasoning (aircraft interaction), (2) Temporal understanding (icing sequence), and (3) Cross-frame reasoning (narrative connection in stick figures), where the baseline fails.
For training, we sample 3,000 videos from ActivityNet Yu et al. (2019), whose diversity of 203 complex activities provides a strong foundation for learning the fundamental reasoning skills our framework targets. This dataset is not part of our evaluation benchmarks, ensuring our results reflect true generalization. We use a large model (InternVL2.5-38B) to generate queries but all aligned and adversarial responses are generated by the smaller target models themselves. This setup ensures models learn to refine their own outputs rather than simply distilling knowledge from a more capable model. Our structured adversarial sampling pipeline initially generates 90,000 preference pairs, which after rigorous filtering (details in Appendix §E.1), are reduced to 7,020 high-quality pairs. We fine-tune models using the SWIFT Zhao et al. (2025) framework for efficient adaptation. All training and evaluations are performed on four NVIDIA L40S GPUs (48GB each), with a maximum input of 32 frames per video to prevent CUDA out-of-memory errors. We employ LoRA Hu et al. (2022) with rank $r=8$ and $\alpha_{LoRA}=8$. For DPO, we set the scaling factor $\lambda$ to 0.1. The overall training loss (Eq.2) combines three components: spatial ($\alpha_{S}=0.4$), temporal ($\beta_{T}=0.4$), and cross-frame ($\gamma_{C}=0.2$) alignment. We apply VideoPASTA to four diverse foundation models: Qwen2.5-VL (7B) (Bai et al., 2025), LLaVA-NeXT-Interleave (7B) (Li et al., 2025a), LLaVA-OneVision (7B) (Li et al., 2024a), and InternVL2.5 (8B) (Chen et al., 2024c). To ensure fair model-agnostic comparisons, we limit training data to 7,020 high-quality preference pairs for all models. This corresponds to the smallest filtered set (from LLaVA-NeXT-Interleave), with larger sets from other models subsampled accordingly to maintain consistency in data quantity. Evaluation is conducted using LMMs-Eval Zhang et al. (2024a) to ensure fair comparisons with prior work.
The Appendix contains additional experiments and information, including: DPO training dynamics (§A), preference learning on smaller (1B-3B) models (§B), adversarial robustness of VideoPASTA (§C), Qwen2.5-VL-specific ablations (§D), full dataset statistics and adversarial samples (§E), qualitative examples (§F), and all prompt templates (§G).
Benchmarks.
We evaluate on general video understanding benchmarks: TempCompass Liu et al. (2024b) (temporal understanding), PerceptionTest Patraucean et al. (2024) (visual perception), NeXTQA Xiao et al. (2021) (compositional reasoning), and MVBench Li et al. (2024b) (multi-task reasoning). For long-form evaluation, we use LongVideoBench Wu et al. (2024) (hour-long videos), MLVU Zhou et al. (2025) (multi-task, 3-minute to 2-hour videos), and VideoMME Fu et al. (2025) (6 visual domains, 30 subfields, 11-second to 1-hour videos).